# A Survey on 3D Gaussian Splatting

### v7

@@ 4.6 Hybrid Representation
Next we showcase various fascinating uses of 3D GS with specially devised structured information. i) Facial Expression Modeling. Considering the challenge of creating high-fidelity 3D head avatars under sparse view conditions, Gaussian Head Avatar [96] introduced controllable 3D Gaussians and an MLP-based deformation field. Concretely, it captured detailed facial expressions and dynamics by optimizing neutral 3D Gaussians alongside the deformation field, thus ensuring both detail fidelity and expression accuracy. ii) Spatiotemporal Modeling. Yang et al[94] proposed to reconstruct dynamic scenes with deformable 3D Gaussians. The deformable 3D Gaussians are learned in a canonical space, coupled with a deformation field (i.e., a spatial MLP) that models the spatial-temporal dynamics. The proposed method also incorporated an annealing smoothing training mechanism to enhance temporal smoothness without additional computational costs. iii) Style Transfer. Saroha et al[107] proposed GS in style, an advanced approach for real-time neural scene stylization. To maintain a cohesive stylized appearance across multiple views without compromising on rendering speed, they used pre-trained 3D Gaussians coupled with a multi-resolution hash grid and a small MLP to produce stylized views. In a nutshell, incorporating structured information can serve as a complementary part for adapting to tasks that are incompatible with the sparsity and disorder of 3D Gaussians.
4.7 New Rendering Algorithm for 3D Gaussians
@@
Recent works [108, 109, 110] explored ray tracing based rendering algorithms as an alternative. For instance, GaussianTracer [108] introduced a new ray tracing implementation for Gaussian primitives, and devised several accelerating strategies according to the uneven density and interleaved nature of Gaussians. EVER [109] deivsed a physically accurate, constant density ellipsoid representation that allows for the exact computation of the volume rendering integral, rather than relying on somewhat satisfactory approximations. This advancement eliminates popping artifacts.
@@
5.15.25.35.45.55.65.7[111], [112], [113], [114], [115], [116], [117],[118], [119], [120], [121], [122], [123], [124][125], [94], [95], [93], [126], [127],[128], [129], [106], [130], [131], [132][133], [134], [135], [136], [137], [138],[139], [140], [141], [142], [90], [143][144], [145], [146], [147][148], [149], [150], [151][152], [153], [154][155], [156], [157][158], [159], [160], [161][44], [162], [163], [164], [165][166], [167], [168], [143], [169][170], [21], [171], [172], [173]
Figure 6: Typical applications benefited from GS (Sec. 5). Some images are borrowed from [132, 135, 154, 160, 146, 166] and redrawn.
5 Application Areas and Tasks
@@
Building on the rapid advancements in 3D GS, a wide range of innovative applications has emerged across multiple domains (Fig. 6) such as robotics (Sec. 5.1), dynamic scene reconstruction and representation (Sec. 5.2), generation and editing (Sec. 5.3), avatar (Sec. 5.4), medical systems (Sec. 5.5), large-scale scene reconstruction (Sec. 5.6), physics (Sec. 5.7), and even other scientific disciplines [174, 175, 24, 176]. Here, we highlight key examples that underscore the transformative impact and potential of 3D GS and offer a more comprehensive collection in Github.
5.1 Robotics
@@
The evolution of scene representation in robotics has been profoundly shaped by the emergence of NeRF, which revolutionized dense mapping and environmental interaction through implicit neural models. However, NeRF’s computational cost poses a critical bottleneck for real-time robotic applications. The shift from implicit to explicit representation not only accelerates optimization but also unlocks direct access to spatial and structural scene data, making 3D GS a transformative tool for robotics. Its ability to balance high-fidelity reconstruction with computational efficiency positions 3D GS as a cornerstone for advancing robotic perception, manipulation, and navigation in dynamic, real-world environments.
@@
The integration of GS into robotic systems has yielded significant advancements across three core domains. In SLAM, GS-based methods [115, 113, 114, 111, 112, 177, 117, 178, 179, 116, 123, 180, 181, 182, 124] excel in real-time dense mapping but face inherent trade-offs. Visual SLAM frameworks, particularly RGB-D variants [112, 114, 178], leverage depth supervision for geometric fidelity but falter in low-texture or motion-degraded environments. RGB-only approaches [113, 115, 183] circumvent depth sensors but grapple with scale ambiguity and drift. Multi-sensor fusion strategies, such as LiDAR integration [177, 182, 159], enhance robustness in unstructured settings at the cost of calibration complexity. Semantic SLAM [123, 117, 116] extends scene understanding through object-level semantics but struggles with scalability due to lighting sensitivity in color-based methods or computational overhead in feature-based methods. 3D GS based manipulation [118, 119, 120, 121, 122] bypasses the need for auxiliary pose estimation in NeRF-based methods, enabling rapid single-stage tasks like grasping in static environments via geometric and semantic attributes encoded in Gaussian properties. Multi-stage manipulation [118, 120], where environmental dynamics demand real-time map updates, requires explicit modeling of dynamic adjustments (e.g., object motions and interactions), material compliance, etc.
@@
5.2 Dynamic Scene Reconstruction
@@
Dynamic scene reconstruction refers to the process of capturing and representing the three-dimensional structure and appearance of a scene that changes over time [184, 185, 186, 187]. This involves creating a digital model that accurately reflects the geometry, motion, and visual aspects of the objects in the scene as they evolve. Dynamic scene reconstruction is crucial in various applications, e.g., VR/AR, 3D animation, and autonomous driving [188, 189, 190].
@@
The key to adapt 3D GS to dynamic scenes is the modeling of temporal dimension which allows for the representation of scenes that change over time. 3D GS based methods [94, 93, 106, 191, 129, 192, 95, 125, 128, 127, 126, 193, 194, 130, 195, 196, 197, 198, 199] for dynamic scene reconstruction can generally be divided into two main categories as discussed in Sec. 4.5 and Sec. 4.6. The first category utilizes additional fields like spatial MLPs or grids to model deformation (Sec. 4.6). For example, Yang et al[94] first proposed deformable 3D Gaussians tailored for dynamic scenes. These 3D Gaussians are learned in a canonical space and can be used to model spatial-temporal deformation with an implicit deformation field (implemented as an MLP). GaGS [132] devised the voxelization of a set of Gaussian distributions, followed by the use of sparse convolutions to extract geometry-aware features, which are then utilized for deformation learning. On the other hand, the second category is based on the idea that scene changes can be encoded into the 3D Gaussian representation with a specially designed rendering process (Sec. 4.5). For instance, Luiten et al[125] introduced dynamic 3D Gaussians to model dynamic scenes by keeping the properties of 3D Gaussians unchanged over time while allowing their positions and orientations to change. Yang et al[93] designed a 4D Gaussian representation, where additional properties are used to represent 4D rotations and spherindrical harmonics, to approximate the spatial-temporal volume of scenes.
@@
While 3D GS advances dynamic scene reconstruction by modeling per-Gaussian deformations, its reliance on fine-grained primitives limits scalability and robustness. Current methods struggle to balance computational efficiency and precision: small-scale reconstructions unify dynamic and static elements but become intractable in large environments, often requiring manual priors to segment regions —– a barrier in unstructured settings. Furthermore, the absence of object-level motion reasoning leads to artifacts and poor generalization over long sequences. Future work might prioritize object-centric frameworks that hierarchically group Gaussians into persistent entities, enabling efficient large-scale reconstruction through inherent motion disentanglement (dynamic vs. static).
5.3 Generation and Editing
@@
Content generation and editing represent two fundamental and inherently interconnected capabilities in modern AI systems. While generation enables the synthesis of novel digital content from scratch or conditional inputs [200, 201, 202], editing provides the crucial ability to refine, adapt, and manipulate existing content with precise control [203]. Together, these capabilities revolutionize creative workflows by combining initial content creation with iterative refinement, enabling applications from professional content production to interactive consumer tools.
@@
Recent advances in generation [204, 133, 134, 205, 206, 207, 208, 209, 210, 211, 212, 213, 214, 135, 215, 216, 217, 136, 218, 219, 220, 221, 222, 223, 224, 225, 137, 226, 138, 227] have led to the emergence of three main approaches. Optimization based methods [204, 133, 134] distill diffusion priors (gradients) to guide 3D model updates with the score functions. While these methods demonstrate impressive fidelity, they face significant computational overhead due to the necessity of comparing multiple viewpoints during the optimization process. Reconstruction based methods [135, 227, 225] reframe the generation problem as a multi-view reconstruction task utilizing pre-trained multi-view diffusion models. Although this approach offers an intuitive and straightforward solution, it grapples with fundamental limitations in maintaining view consistency. The lack of strict geometric constraints across different viewpoints often results in inconsistent surface geometry and degraded texture quality, particularly in regions with complex visual features. Direct 3D generation methods train diffusion models on 3D representations [138, 220, 226]. While the learned 3D diffusion models facilitate multi-view consistency, the demanding computational costs impede the expansion of training scales necessary for improved generative diversity.
@@
Current editing works [141, 142, 228, 91, 92, 90, 229, 128, 127, 126, 230, 231, 232, 233, 234, 235, 236, 143, 237, 140, 238, 239] fall into two primary classes. The first class leverages 2D image-editing models (e.g., diffusion-based editors) to iteratively refine 3D Gaussians. Early efforts [141, 142, 233] adopt optimization- or reconstruction-based strategies akin to methods in generation, but introduce task-specific control signals. However, naively applying 2D edits independently across views often introduces multi-view inconsistencies. Subsequent works [240, 140, 238, 239] mitigate this through iterative refinement or cross-view attention, albeit at increased computational costs for alignment. A notable challenge is unintended object deformations, attributed to the weak 3D geometric priors in 2D editing models and the difficulty of reconciling 2D edits with underlying 3D structures. The second class exploits the explicit nature of 3D GS to enable direct manipulation based on embedded properties such as semantics [232, 92, 91, 143] and key points [128]. However, this class remains underexplored due to essentail challenges: the lack of inherent ordering of Gaussians complicates the design of efficient indexing schemes, while editing attributes (e.g., texture and geometry) requires careful regularization and alignment to preserve plausibility.
5.4 Avatar

### v8

@@ 4.6. Hybrid Representation
Next we showcase various fascinating uses of 3D GS with specially devised structured information. i) Facial Expression Modeling. Considering the challenge of creating high-fidelity 3D head avatars under sparse view conditions, Gaussian Head Avatar (Xu et al., 2024a) introduced controllable 3D Gaussians and an MLP-based deformation field. Concretely, it captured detailed facial expressions and dynamics by optimizing neutral 3D Gaussians alongside the deformation field, thus ensuring both detail fidelity and expression accuracy. ii) Spatiotemporal Modeling. Yang et al. (Yang et al., 2024b) proposed to reconstruct dynamic scenes with deformable 3D Gaussians. The deformable 3D Gaussians are learned in a canonical space, coupled with a deformation field (i.e., a spatial MLP) that models the spatial-temporal dynamics. The proposed method also incorporated an annealing smoothing training mechanism to enhance temporal smoothness without additional computational costs. iii) Style Transfer. Saroha et al. (Saroha et al., 2024) proposed GS in style, an advanced approach for real-time neural scene stylization. To maintain a cohesive stylized appearance across multiple views without compromising on rendering speed, they used pre-trained 3D Gaussians coupled with a multi-resolution hash grid and a small MLP to produce stylized views. In a nutshell, incorporating structured information can serve as a complementary part for adapting to tasks that are incompatible with the sparsity and disorder of 3D Gaussians.
4.7. New Rendering Algorithm for 3D Gaussians
@@
Recent works (Moenne-Loccoz et al., 2024; Mai et al., 2024; Condor et al., 2025) explored ray tracing based rendering algorithms as an alternative. For instance, GaussianTracer (Moenne-Loccoz et al., 2024) introduced a new ray tracing implementation for Gaussian primitives, and devised several accelerating strategies according to the uneven density and interleaved nature of Gaussians. EVER (Mai et al., 2024) deivsed a physically accurate, constant density ellipsoid representation that allows for the exact computation of the volume rendering integral, rather than relying on somewhat satisfactory approximations. This advancement eliminates popping artifacts.
@@
5. Application Areas and Tasks
@@
Building on the rapid advancements in 3D GS, a wide range of innovative applications has emerged across multiple domains (c.f. Fig. 6) such as robotics (Sec. 5.1), dynamic scene reconstruction and representation (Sec. 5.2), generation and editing (Sec. 5.3), avatar (Sec. 5.4), medical systems (Sec. 5.5), large-scale scene reconstruction (Sec. 5.6), physics (Sec. 5.7), and even other scientific disciplines (Zhang et al., 2024j; Wu et al., 2024f; Zhang et al., 2024e; Li et al., 2024d). Here, we highlight key examples that underscore the transformative impact and potential of 3D GS and offer a more comprehensive collection in Github. We also provide a collection of representative datasets (c.f. Table 5.7) according to our taxonomy.
5.15.25.35.45.55.65.7(Yan et al., 2024c), (Keetha et al., 2024), (Matsuki et al., 2024), (Yugay et al., 2023), (Huang et al., 2024c), (Li et al., 2024e), (Ji et al., 2024a)(Lu et al., 2024c), (Abou-Chakra et al., 2024), (Shorinwa et al., 2024), (Ji et al., 2024b), (Zheng et al., 2024a), (Zhu et al., 2024b), (Peng et al., 2024a)(Luiten et al., 2024), (Yang et al., 2024b), (Wu et al., 2024c), (Yang et al., 2024d), (Yu et al., 2024b), (Shao et al., 2024b)(Huang et al., 2024d), (Das et al., 2024), (Lin et al., 2024a), (Li et al., 2024a), (Sun et al., 2024a), (Lu et al., 2024a)(Tang et al., 2024b), (Yi et al., 2024), (Tang et al., 2024a), (Zhou et al., 2024b), (Li et al., 2024b), (Mu et al., 2024)(Jiang et al., 2024a), (Wang et al., 2024a), (Chen et al., 2024a), (Fang et al., 2024), (Zhou et al., 2024a), (Qiu et al., 2024)(Li et al., 2024k), (Hu and Liu, 2024), (Lei et al., 2024a), (Yuan et al., 2024)(Zhou et al., 2024d), (Qian et al., 2024a), (Dhamo et al., 2023), (Li et al., 2024j)(Huang et al., 2024b), (Liu et al., 2024b), (Zhu et al., 2024c)(Zhao et al., 2024d), (Wang et al., 2024c), (Bonilla et al., 2024)(Wu et al., 2024g), (Wu et al., 2024b), (Xiong et al., 2024), (Zhao et al., 2024c)(Cheng et al., 2024), (Kerbl et al., 2024), (Liu et al., 2024a), (Lin et al., 2024b), (Ren et al., 2024)(Xie et al., 2024), (Liu et al., 2024e), (Borycki et al., 2024), (Qiu et al., 2024), (Huang et al., 2025)(Zhang et al., 2024h), (Jiang et al., 2024c), (Feng et al., 2024a), (Zhong et al., 2024), (Shao et al., 2024a)
Figure 6. Typical applications benefited from GS (Sec. 5). Some images are borrowed from (Lu et al., 2024a; Tang et al., 2024a; Zhu et al., 2024c; Xiong et al., 2024; Lei et al., 2024a; Xie et al., 2024) and redrawn.
5.1. Robotics
@@
The evolution of scene representation in robotics has been profoundly shaped by NeRF. However, NeRF’s computational cost poses a critical bottleneck for real-time applications. The shift from implicit to explicit representation not only accelerates optimization but also unlocks direct access to spatial and structural scene data, making 3D GS a transformative tool for robotics. Its ability to balance high-fidelity reconstruction with computational efficiency positions 3D GS as a cornerstone for advancing robotic perception, manipulation, and navigation in real-world environments.
@@
The integration of GS into robotic systems has yielded significant advancements across three core domains. In SLAM, GS-based methods (Huang et al., 2024c; Matsuki et al., 2024; Yugay et al., 2023; Yan et al., 2024c; Keetha et al., 2024; Hong et al., 2024; Ji et al., 2024a; Sun et al., 2024b; Tosi et al., 2024; Li et al., 2024e; Zhu et al., 2024b; Deng et al., 2024; Hu et al., 2024a; Lang et al., 2024; Peng et al., 2024a) excel in real-time dense mapping but face inherent trade-offs. Visual SLAM frameworks, particularly RGB-D variants (Keetha et al., 2024; Yugay et al., 2023; Sun et al., 2024b), leverage depth supervision for geometric fidelity but falter in low-texture or motion-degraded environments. RGB-only approaches (Matsuki et al., 2024; Huang et al., 2024c; Sandström et al., 2024) circumvent depth sensors but grapple with scale ambiguity and drift. Multi-sensor fusion strategies, such as LiDAR integration (Hong et al., 2024; Lang et al., 2024; Wu et al., 2024b), enhance robustness in unstructured settings at the cost of calibration complexity. Semantic SLAM (Zhu et al., 2024b; Ji et al., 2024a; Li et al., 2024e) extends scene understanding through object-level semantics but struggles with scalability due to lighting sensitivity in color-based methods or computational overhead in feature-based methods. 3D GS based manipulation (Lu et al., 2024c; Abou-Chakra et al., 2024; Shorinwa et al., 2024; Ji et al., 2024b; Zheng et al., 2024a) bypasses the need for auxiliary pose estimation in NeRF-based methods, enabling rapid single-stage tasks like grasping in static environments via geometric and semantic attributes encoded in Gaussian properties. Multi-stage manipulation (Lu et al., 2024c; Shorinwa et al., 2024), where environmental dynamics demand real-time map updates, requires explicit modeling of dynamic adjustments (e.g., object motions and interactions), material compliance, etc.
@@
5.2. Dynamic Scene Reconstruction
@@
Dynamic scene reconstruction refers to the process of capturing and representing the three-dimensional structure and appearance of a scene that changes over time (Pumarola et al., 2021; Park et al., 2021b, a; Guo et al., 2023). This involves creating a digital model that accurately reflects the geometry, motion, and visual aspects of the objects in the scene as they evolve. Dynamic scene reconstruction is crucial in various applications, e.g., VR/AR, 3D animation, and autonomous driving (Zhou et al., 2024c; Yan et al., 2024a; Zhou et al., 2024e).
@@
The key to adapt 3D GS to dynamic scenes is the modeling of temporal dimension which allows for the representation of scenes that change over time. 3D GS based methods (Yang et al., 2024b, d; Lin et al., 2024a; Kratimenos et al., 2023; Das et al., 2024; Shaw et al., 2023; Wu et al., 2024c; Luiten et al., 2024; Huang et al., 2024d; Shao et al., 2024b; Yu et al., 2024b; Liang et al., 2023; Katsumata et al., 2023; Li et al., 2024a; Guo et al., 2024; Bae et al., 2024; Lei et al., 2024b; Wang et al., 2024d; Duan et al., 2024) for dynamic scene reconstruction can generally be divided into two main categories as discussed in Sec. 4.5 and Sec. 4.6. The first category utilizes additional fields like spatial MLPs or grids to model deformation (Sec. 4.6). For example, Yang et al. (Yang et al., 2024b) first proposed deformable 3D Gaussians tailored for dynamic scenes. These 3D Gaussians are learned in a canonical space and can be used to model spatial-temporal deformation with an implicit deformation field (implemented as an MLP). GaGS (Lu et al., 2024a) devised the voxelization of a set of Gaussian distributions, followed by the use of sparse convolutions to extract geometry-aware features, which are then utilized for deformation learning. On the other hand, the second category is based on the idea that scene changes can be encoded into the 3D Gaussian representation with a specially designed rendering process (Sec. 4.5). For instance, Luiten et al. (Luiten et al., 2024) introduced dynamic 3D Gaussians to model dynamic scenes by keeping the properties of 3D Gaussians unchanged over time while allowing their positions and orientations to change. Yang et al. (Yang et al., 2024d) designed a 4D Gaussian representation, where additional properties are used to represent 4D rotations and spherindrical harmonics, to approximate the spatial-temporal volume of scenes.
@@
While 3D GS advances dynamic scene reconstruction by modeling per-Gaussian deformations, its reliance on fine-grained primitives limits scalability and robustness. Current methods struggle to balance computational efficiency and precision: small-scale reconstructions unify dynamic and static elements but become intractable in large environments, often requiring manual priors to segment regions —– a barrier in unstructured settings. Furthermore, the absence of object-level motion reasoning leads to artifacts and poor generalization over long sequences. Future work might prioritize object-centric frameworks that group Gaussians into persistent entities so as to model inherent motion disentanglement (dynamic vs. static).
5.3. Generation and Editing
@@
Content generation and editing represent two fundamental and inherently interconnected capabilities in modern AI systems. While generation enables the synthesis of novel digital content from scratch or conditional inputs (Goodfellow et al., 2020; Ho et al., 2020; Rombach et al., 2022), editing provides the crucial ability to refine, adapt, and manipulate existing content with precise control (Zhang et al., 2023a). Together, these capabilities revolutionize creative workflows by combining initial content creation with iterative refinement, enabling applications from professional content production to interactive consumer tools.
@@
Recent advances in generation (Chen et al., 2024b; Tang et al., 2024b; Yi et al., 2024; Liang et al., 2024a; Liu et al., 2024f; Yang et al., 2023a; Zou et al., 2024; Ling et al., 2024; Ren et al., 2023; Yin et al., 2023; Zhang et al., 2023b; Pan et al., 2024; Xu et al., 2024c; Tang et al., 2024a; Yang et al., 2024c; Barthel et al., 2024; Jiang and Wang, 2024; Zhou et al., 2024b; Zhuo et al., 2024; Wu et al., 2024d; He et al., 2024; Yang and Wang, 2024; Kim et al., 2024; Feng et al., 2024b; Li et al., 2024g; Melas-Kyriazi et al., 2024; Li et al., 2024b; Zhang et al., 2024b; Mu et al., 2024; Lee et al., 2024a) have led to the emergence of three main approaches. Optimization based methods (Chen et al., 2024b; Tang et al., 2024b; Yi et al., 2024) distill diffusion priors (gradients) to guide 3D model updates with the score functions. While these methods demonstrate impressive fidelity, they face significant computational overhead due to the necessity of comparing multiple viewpoints during the optimization process. Reconstruction based methods (Tang et al., 2024a; Lee et al., 2024a; Melas-Kyriazi et al., 2024) reframe the generation problem as a multi-view reconstruction task utilizing pre-trained multi-view diffusion models. Although this approach offers an intuitive and straightforward solution, it grapples with fundamental limitations in maintaining view consistency. The lack of strict geometric constraints across different viewpoints often results in inconsistent surface geometry and degraded texture quality, particularly in regions with complex visual features. Direct 3D generation methods train diffusion models on 3D representations (Mu et al., 2024; He et al., 2024; Zhang et al., 2024b). While the learned 3D diffusion models facilitate multi-view consistency, the demanding computational costs impede the expansion of training scales necessary for improved generative diversity.
@@
Current editing works (Chen et al., 2024a; Fang et al., 2024; Huang and Yu, 2023; Ye et al., 2024; Cen et al., 2023; Zhou et al., 2024a; Lan et al., 2023; Huang et al., 2024d; Shao et al., 2024b; Yu et al., 2024b; Zhuang et al., 2024; Dou et al., 2024; Hu et al., 2024b; Palandra et al., 2024; Gu et al., 2024; Lyu et al., 2024; Liu et al., 2024c; Qiu et al., 2024; Zhang et al., 2024a; Wang et al., 2024a; Zhang et al., 2024g; Wu et al., 2024a) fall into two primary classes. The first class leverages 2D image-editing models (e.g., diffusion-based editors) to iteratively refine 3D Gaussians. Early efforts (Chen et al., 2024a; Fang et al., 2024; Palandra et al., 2024) adopt optimization- or reconstruction-based strategies akin to methods in generation, but introduce task-specific control signals. However, naively applying 2D edits independently across views often introduces multi-view inconsistencies. Subsequent works (Wang et al., 2024e, a; Zhang et al., 2024g; Wu et al., 2024a) mitigate this through iterative refinement or cross-view attention, albeit at increased computational costs for alignment. A notable challenge is unintended object deformations, attributed to the weak 3D geometric priors in 2D editing models and the difficulty of reconciling 2D edits with underlying 3D structures. The second class exploits the explicit nature of 3D GS to enable direct manipulation based on embedded properties such as semantics (Hu et al., 2024b; Cen et al., 2023; Ye et al., 2024; Qiu et al., 2024) and key points (Huang et al., 2024d). However, this class remains underexplored due to essential challenges: the lack of inherent ordering complicates the design of efficient indexing schemes, while editing attributes (e.g., texture and geometry) requires careful regularization and alignment to preserve plausibility.
5.4. Avatar